A circular railway track of radius 400m is to be made so that a train can move with a speed of 10m/s without friction. If the distance between the two rails is 1m by how much the outer rail should be raised? (g=10m/s2)

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
tan theta = v²/ rg = 100/4000 = 1/40
Now you can solve this.
 
Regards
Arun (askIITians forum expert)